In 1990, Michael's college tuition was $1500. In 1995, tuition at the same college cost $3200. Assuming that an exponential model applies, how much will tuition cost in 2009?? I think the answer is $6837 ????????
Answer by checkley71(8403)

(1905-1990)/(3,000-1500)
5/1500=(2009-1990)/X-1500)
5/1500=19/(X-1500)
5(X-1500)=19*1500
5X-7500=28500
5X=28500+7500
5X=36000
X=36000/5
X=$7200 IS THE PROJECTED TUITION IN 2009
PROOF:
5/1500=19/(7200-1500)
5/1500=19/5700
19*1500=5*5700
28500=28500